OpenStudy (anonymous):

Which exponential function goes through the points (1, 8) and (4, 512)? f(x) = 4(2)x f(x) = 8(2)-x f(x) = 4(4)-x f(x) = 2(4)x

OpenStudy (igreen):

The last one. You can graph them here: https://www.desmos.com/calculator

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

The function f(x) is shown below: f(x) = (4)x + 5 The function f(x) is shifted to the right by 13 units. Which of the following best represents the new function? f(x) = (4)x - 8 f(x) = (4)x + 13 f(x) = (4)x + 18 f(x) = (4)x - 13

OpenStudy (anonymous):

Sabrina is using an online calculator to calculate the outputs f(n) for different inputs n. The ordered pairs below show Sabrina's inputs and the corresponding outputs displayed by the calculator: (1, 6), (2, 11), (3, 16), (4, 21) Which of the following functions best represents the rule that the calculator uses to display the outputs? f(n) = 5n - 1 f(n) = 5n + 1 f(n) = 4n + 3 f(n) = 4n - 3

OpenStudy (anonymous):

Wendy initially has 6 hours of pop music and 3 hours of classical music in her collection. Every month onwards, the hours of pop music in her collection is 9% more than what she had the previous month. Her classical music does not change. Which function shows the total hours of music she will have in her collection after x months? f(x) = 6(1.09)x + 3 f(x) = 3(1.09)x + 6 f(x) = 6(0.09)x + 3 f(x) = 3(0.09)x+ 6

OpenStudy (anonymous):

The function f(x) = (1.006235)12x models the monthly interest that a bank offers to Dan after x years. Dan converts the function to have x isolated in the exponent. What is the approximate rate of growth? 1% 8% 10% 12%

OpenStudy (anonymous):

The function below shows the number of students of a school who enrolled for ballet classes. Let f(x) represent the total number of students who enrolled for the classes after x years: f(x) = 12(1.06)x The average rate of change in the number of students who enrolled for ballet classes from the third to the fifth year is ________students per year. Round your answer to the nearest whole number.

OpenStudy (anonymous):

thats what i need help with

OpenStudy (igreen):

That's a lot of questions..

OpenStudy (igreen):

I think you got the question to #1 right..

OpenStudy (anonymous):

ok what about all the other ones

OpenStudy (igreen):

The function f(x) is shown below: f(x) = (4)x + 5 The function f(x) is shifted to the right by 13 units. Which of the following best represents the new function? f(x) = (4)x - 8 f(x) = (4)x + 13 f(x) = (4)x + 18 f(x) = (4)x - 13 <-- Answer

OpenStudy (muzzack):

yes your answer is correct

OpenStudy (igreen):

Sabrina is using an online calculator to calculate the outputs f(n) for different inputs n. The ordered pairs below show Sabrina's inputs and the corresponding outputs displayed by the calculator: (1, 6), (2, 11), (3, 16), (4, 21) Which of the following functions best represents the rule that the calculator uses to display the outputs? f(n) = 5n - 1 f(n) = 5n + 1 <-- Answer f(n) = 4n + 3 f(n) = 4n - 3

OpenStudy (igreen):

I can't do the radicals one's..They're too difficult..

OpenStudy (igreen):

Wendy initially has 6 hours of pop music and 3 hours of classical music in her collection. Every month onwards, the hours of pop music in her collection is 9% more than what she had the previous month. Her classical music does not change. Which function shows the total hours of music she will have in her collection after x months? f(x) = 6(1.09)x + 3 <-- Answer f(x) = 3(1.09)x + 6 f(x) = 6(0.09)x + 3 f(x) = 3(0.09)x+ 6

OpenStudy (igreen):

The function f(x) = (1.006235)^12x models the monthly interest that a bank offers to Dan after x years. Dan converts the function to have x isolated in the exponent. What is the approximate rate of growth? 1% 8% <-- Answer 10% 12%

OpenStudy (igreen):

Is the 'x' an exponent in this? f(x) = 12(1.06)x

OpenStudy (igreen):

@1111111

OpenStudy (igreen):

You can give me a medal by clicking 'Best Response'. @1111111

OpenStudy (igreen):

Which one is that?

OpenStudy (anonymous):

@iGreen this is number 5

OpenStudy (igreen):

The answer to that one is C.

OpenStudy (anonymous):

AND WHAT ABOUT ALL THE OTHER ONES

OpenStudy (anonymous):

oops sorry i hit the caps locks by accident

OpenStudy (igreen):

Does "f(x) = 12(1.06)x" look like this? \(f(x) = 12(1.06)^x\) or like this? \(f(x) = 12 (1.06)x\)

OpenStudy (anonymous):

f(x)=12(1.06)^x

OpenStudy (igreen):

Okay

OpenStudy (igreen):

(3, 14.292) (5, 16.059) \(m = \Large\frac{y2-y1}{x2-x1}\) \(m = \Large\frac{16.059-14.292}{5-3}\) \(m = \Large\frac{1.76700}{2}\) \(m = 0.8835\)

OpenStudy (igreen):

Round to the nearest whole number = 1.

OpenStudy (anonymous):

thanks and what about the last two questions

OpenStudy (igreen):

And I don't know the last one..sorry.

OpenStudy (anonymous):

ok thats fine

OpenStudy (anonymous):

and what about hte one where iut says evaluate 5 blah blah blah

OpenStudy (anonymous):

@iGreen what about this one

OpenStudy (anonymous):

i dont get this one

OpenStudy (igreen):

I don't know it either..